If a system has 4.50 x 102 kcal of work done to it, and releases 5.00 × 102 kJ of heat into its surroundings. what is the change in internal energy of the system? Number kJ
Answer #1

Using the first law of thermodynamics

∆U = ∆q+∆W

∆Q = - 5.00×102kJ ( negative sign indicate that heat released by system)

∆W =+ 4.50×102kJ ( positive sign indicate that work done on the system by the surrounding)

Putting both value in formula

∆U = -2.00×102kJ + 4.5×102 kJ = 2.5×102kJ